### v3

# Pixel column issue in the ATLAS Inner Tracker modules

Abstract
Pixel modules are currently being built for the ATLAS ITk Pixel detector upgrade. During the preproduction phase, recurring chip malfunctioning was observed during electrical testing. It was possible to bypass this issue by disabling some pixel core columns in the ITkPix readout chip. Therefore the issue is called “core column issue” which is a direct disqualifier for a pixel module. A concerning number of cases has been observed in pixel modules with ITkPix v1.1 as well as v2 chips which significantly impacts the module yield. However, the behaviour is erratic and there is not any evidence hinting at the origin of this issue. These proceedings outline the investigations of the issue, highlighting the electrical behaviour during testing, present findings from the data collected via our production database and through visual inspection, and point towards possible causes of the issue.
1 Introduction
The Inner Tracker (ITk) is the upgrade tracker for ATLAS [1] in the High Luminosity (HL)- Large Hadron Collider (LHC) era. The pixel detector [2] is currently in its production phase, with more than 10000 hybrid modules being built. The modules come in two flavours: quad modules (figure 1(a)) consisting of four front-end (FE) chips and a quad-size planar sensor tile of about $4\times 4$ cm2, and triplet modules of different geometries, figure 1(b), consisting of three bare modules, each made of one FE chip and one single 3D sensor tile. In the ITk pixel detector, about 95% of the modules are quad modules while 5% are triplet modules.
The FE chip was developed by the RD53 collaboration [3] for both the ATLAS and CMS inner tracker upgrade for the HL-LHC. They differ only in the size of the matrix and in the analog pixel circuit. The ATLAS variant is called "ITkPix" and consists of 400 columns $\times$ 384 rows, resulting in a total of 153600 pixels. The digital logic is grouped into 50 core columns as indicated in figure 2(a), where each core column consists of 8 pixel columns.
Figure 1: Two pixel module flavours: (a) a quad module, (b) a linear triplet module.
Figure 2: (a) The ITkPix chip layout: pixel matrix (left) with core column logic (right). (b) Occupancy map from a digital scan with 100 injections on a module with some core columns disabled (marked in green). The bottom right chip was not wirebonded.
2 Core Column Issue
Throughout module preproduction with ITkPix v1.1 and during production using ITkPix v2, certain chip malfunctioning is observed on chips classified as fully functionining - "green" - by wafer probing. The symptoms are erratic and not consistent, e.g. chip readout is stuck, or chip sends continuous bad data etc. Chip operation can be recovered by disabling one or more core columns (figure 2(b)) so that the chip behaves normally. Thus this kind of chip malfunctioning is called "core column issue". Because a core column contains 2% of the pixels on the chip, disabling one core column would directly fail a module by our current quality control (QC) criteria (total $<$1% bad pixels). This issue is observed on ITkPix v1.1 modules and ITkPix v2 modules, as well as by CMS on their CMS Readout Chip (CROC) v1 modules.
2.1 Statistics
To obtain an overview of the frequency with which the core column issue occurs, queries were made to the ITk production database (PDB) [4] which records information of all components, their stages and tests. FE chip configurations were queried to obtain the settings of the core-column-enable registers which contain information about disabled core columns. A snapshot of the CC statistics is shown in table 1. The data are sorted by chip version and testing temperature in order to show any systematics.
Table 1: Data extracted from the production database. The columns "CC/chips(modules)" are fractions of chip (modules) with at least one disabled CCs out of the total number of chips (modules). The testing temperatures "warm" and "cold" correspond to about 20 C and $-15\,^{\circ}$, respectively.
chip version testing temperature CC/modules CC/modules (%) CC/chips CC/chips (%)
v1.1 warm 77/357 21.6 116/1393 8.3
cold 42/249 16.9 57/982 5.8
v2 warm 17/118 14.4 25/470 5.3
cold 9/85 10.6 12/340 3.5
The largest amount of the data are from warm tests on v1.1 chips assembled in both quad and triplet modules, therefore they are used in the following investigations. Figure 3 shows fractions of chips (3(a)) or modules (3(b)) with CC issues, which are similar across all hybridisation vendors, showing that the issues are not associated with one vendor. The number of FE chips that are disabled in software during testing is also indicated in figure 3(a). This could be due to either the user not associating errors with the core column issue, or other issues such as the chip is not wirebonded due to damage of the wirebond pads or wirebonding failure. Therefore, disabled chips are not included in the number of FE chips with CC.
Figure 3: Fraction of chips (a) and modules (b) with disabled core columns per hybridisation vendor. (c) positions of disabled CCs on all problematic chips.
Figure 3(c) shows the distribution of all disabled CCs for all hybridisation vendors. The distribution is uniform across FE chip except for the edges where an increase is seen. The increase in disabled edge columns is associated with one vendor. The position of chips on wafers were studied as shown in figure 4. While the cumulative map (figure 4(a)) does not show any preferencial position for CC chips, looking at individual wafer maps reveals that CC chips often cluster together in a row or in a column which could indicate problems with dicing or picking, like shown in figure 4(b).
Other variables, such as the assembly or testing date or site, sensor leakage current etc, showed no relation to the core column issue.
Figure 4: (a) Cumulative wafer map with occurence of CC chip position. (b) example wafer map where CC chips often appear in consecutive oder in a row or column. The z-scale indicate the location of disabled core columns on that chip.
3 Investigations
Intrisic issues of the design or the chip itself can be excluded to be origins of the core column issue because such would have been identified, flagged or rejected accordingly during the probing of the FE chip wafers. The only stages at which changes happen to the chips are between wafer probing and initial testing, which would be hybridisation and module assembly. In the hybridisation stage, chips are bump-deposited, thinned, diced and flip-chipped into a bare module. The bare module is glued onto a flexible module printed circuit board (PCB) and wirebonded in the assembly stage. Especially hybridisation is a delicate process during which mechanical damages can occur.
3.1 Probing of diced chips
To test this hypothesis, 63 and 25 diced ITkPix v2 chips from two different wafers were probed. While the former were all probed to be "green" on the wafer level, the latter were not probed prior to dicing. Only four from the originally green chips had any pixel issues, which are still categorised as "green" after probing. Two of them had bad double-column bias to the analog FE circuit, and the other two chips likely had data acquisition (DAQ) instability that showed up in the scan mask pattern. None of them showed CC issues based on what has been observed.
3.2 Inconsistencies and analog dependencies
Bad core columns often manifest themselves already at register reading or in digital scans, but also analog dependencies have been observed.
Very few CC cases can be potentially narrowed down to bad pixel regions (4$\times$1 pixels) as shown in figure 5(a) or bad pixel cores (8$\times$8 pixels) as shown in figure 5(b). Once these smaller areas are disabled, the chip can function normally without disabling the CCs. This would drastically reduce the number of bad pixels overall on a chip compared to disabling entire core columns.
CC issue often shows inconsistent symptoms even within the same chip. One example is figure 5(b), where the bad pixel core was hard to find. While the software diagnostics tool cannot find any bad CCs in this chip, the analog scan often fails, indicating CC issues. Only when sometimes an analog scan runs successfully, this bad pixel core can be seen in the occupancy map.
CC symptoms sometimes show dependencies on the analog settings, for example threshold or injection charge. Figure 5(c) shows a bad occupancy map of the chip with CC issue when the injected charge is 1650 e. The same chip behaves normally when the injected charge is 2600 e as shown in figure 5(d).
When a chip gets stuck due to bad CCs, a reset signal must be issued, which is not the normal mode of operation. Modules with continuous data output were studied to confirm that these data cannot be decoded.
Figure 5: (a) Map from a pixel test highlighting a bad 4$\times$1 pixel region. (b) Occupancy map from an analog scan with a bad 8$\times$8 pixel core with mostly 0 hits. (c) Occupancy map of a module at 1650 e injection charge. (d) Occupancy map of the same module at 2600 e injection charge.
3.3 Mechanical damage
Core column issues were observed on badly diced chips. Therefore, we tried to obtain a correlation with mechanical damage.
First hint
[…]
Figure 6: A quad module has developed core column issue. Careful visual inspection revealed a crack within the area of disabled core columns which came from a solder splash on the vacuum chuck.
Visual inspection
Visual inspection of many affected modules were done. While a few known chips with bad dicing edges, as shown in figure 7(a), had core column issues, most CC modules collected for this campaign did not show any vislble bad dicing as shown in figure 7(b), or any damage on accessible part of the chip, e.g. the backside or around the wirebond pads. There was no clear correlation between modules with CC issue due to bad dicing. Bad CC issues can be associated with poor dicing, but poor dicing does not necessarily imply CC issue.
Figure 7: Different quality of dicing edges in v1.1 and v2 chips. Core column issues have been observed both on chips with bad (a) and good (b) dicing edges.
Induce surface damage
From the diced single ITkPix v2 chips in section 3.1, mechanical damage was induced trying to provoke core column issue. Each chip is divided into several sections as shown in figure 8(a). In each of the sections, a different scratch pattern with varying strength was induced via probe needles. An example of a scratch is shown in figure 8(b). The chips were then probed again to test for any induced damage. From the two scratched chips, one ended up dead, the other one suffered a damaged double-column analog bias from one of the scratches, as can be seen in an analog scan in figure 8(c). This agrees with expectations as the power distribution of the chip happens in the top metal layers. However, core column issues were not induced by this scratching test.
Figure 8: Destructive test on diced v2 chips by inducing different scratch patterns in different regions of a chip (a) and (b). The result was a chip with partially damaged double-column analog bias that is visible in an analog scan (c).
Rejected chips
Chip defects were studied that were grounds for rejection by hybridisation vendors. Six vendor-rejected chips with different defects were collected, visually inspected with a higher magnification and tested on a probe station. Using the same software to test all these chips, the two chips with what looks like chipped edges as shown in the top row in figure 9 had CC issues while no core column issues were found on the other four chips with bump defects as shown in the bottom row in figure 9.
Figure 9: Top: damages of two vendor-rejected chips which exhibit core column issue. Bottom: defects of four vendor-rejected chips without core column issue.
Destructive tests
Two modules with CC issues from the same vendor were disassembled by heating up on a hotplate, shown in figure 10(a), and using vacuum tools to pull apart the chips at the melted solder bumps. Three defects were found on one of the chips, shown as an example in figure 10(b), where defect 1 and 3 are similar as can be seen in figure 11. The location of disabled CCs agrees with the location of defect 3. The appearence of the defect correlated with the CC issue on the second module is the same as defect 3. An energy-dispersive x-ray (EDX) measurement revealed that the extra material consists mostly of silicon, which may be due to some debris from dicing that got between the FE chip and the sensor during hybridisation.
Figure 10: (a) A module on a hotplate (left). After successful disassembly, four chips can be seen on the hotplate (right). (b) Defects found on one disassembled chip. Out of the three defects, the location of defect 3 agrees with the location of the disabled CCs in a scan (c).
Figure 11: Top: defect 1 does not correspond to a disabled CC. Bottom: defect 3 is associated with a disabled CC (same appearance as on the second disassembled CC module). (a) birds-eye view of the defect. (b) 3D view of the defect. (c) measured dimensions of the defect.
4 Conclusion
The core column issue observed on pixel modules affects about 20–25% of ITkPix v1.1 and v2 modules. Our investigations uncovered inconsistent symptoms in the behaviour of the chip, showed potential dependency with chip’s analog settings and confirmed the cause being mechanical damage as we first suspected. While we are investigating deployment of enhanced QC procedures at hybridisation vendors, we also prepare to check for bad core columns within our module QC procedure by using the software tools developed. However, due to the origin of the issue, the development throughout the whole QC process and longterm impact are not yet known, especially regarding handling, thermal stress and irradiation. Impact on tracking performance is being studied via simulations.